an experise nurse can do a job in 20 minutes..an unexperise nurse and a experise nurse can do a job in 12 minutes...what did the unexperise nurse do the job in...do u use w=r times t...explain please...
Answer by Alan3354(69443)

an experise nurse can do a job in 20 minutes..an unexperise nurse and a experise nurse can do a job in 12 minutes...what did the unexperise nurse do the job in...do u use w=r times t...explain please...
--------------------
I think you mean experienced and inexperienced.
Anyway, call the experienced x, and the other y to save typing.
-------------
x does the job in 20 minutes, so she does 1/20 of the job per minute.
y does the job in t minutes, so the does 1/t of the job per minute.
Together, they do 1/20 + 1/t = 1/12
Solve for t.
1/t = 1/12 - 1/20 = 10/120 - 6/120 = 4/120
t = 120/4 = 30 minutes.
The inexperienced nurse takes 30 minutes.